# Imperative Learning: A Self-supervised Neuro-Symbolic Learning Framework for Robot Autonomy

### v6

@@ 5.1 Path Planning
In this section, we will conduct the experiments for iA and iPlanner for global and local planning, respectively.
@@
To demonstrate the effectiveness of IL in global planning, we conduct both qualitative and quantitative evaluations for iA.
@@ 5.1.1 Baselines
To evaluate the generalization ability of the proposed iA framework, we select both classic and data-driven methods as the baselines. Specifically, we will compare iA with the popular classic method, weighted A (Pohl, 1970) and the latest imitation learning-based methods including neural A (Yonetani et al., 2021b), focal search with path probability map (FS + PPM) and weighted A with correction factor (WA + CF) (Kirilenko et al., 2023).
@@ 5.1.2 Datasets
To evaluate iA in diverse environments, we select three widely used datasets in path planning, Motion Planning (MP) (Bhardwaj et al., 2017), Matterport3D (Chang et al., 2017), and Maze (Ivanitskiy et al., 2023). Specifically, the MP dataset contains a set of binary obstacle maps from eight environments; the Matterport3D dataset provides diverse indoor truncated signed distance function (TSDF) maps; and the Maze dataset contains sophisticated mazes-type maps. To better evaluate the generalization ability, we randomly sample the start and goal positions following (Yonetani et al., 2021b) to provide more diverse maps. It is essential to note that all methods only use the MP dataset for training, and use Matterport3D and Maze for testing, which is to demonstrate their generalization abilities.
5.1.3 Metrics
To evaluate the performance of different algorithms, we adopt the widely used metrics including the reduction ratio of node explorations (Exp) (Yonetani et al., 2021b) and the reduction ratio of time (Rt). Exp is the ratio of the reduced search area relative to the classical A, which can be defined as $\textit{Exp}=100\times\frac{S^{*}-S}{S^{*}}$, where $S$ represents the search area of the method and $S^{*}$ is the search area of classical A. Rt is the ratio of saved time relative to the classical A, which can be defined as $\textit{Rt}=100\times\frac{t^{*}-t}{t^{*}}$, where $t$ is the runtime of the method and $t^{*}$ is the runtime of classical A. Intuitively, Exp aims to measure the search efficiency while Rt aims to measure the runtime efficiency.
5.1.4 Quantitative Performance
Table 2: Path planning evaluation on the MP dataset. The symbol ‘--’ denotes the method cannot provide meaningful results with that map size due to the large search space.
@@ 5.1.4 Quantitative Performance
WA+CF 47.3% 45.3% -- -- -- --
FS+PPM --8.9% --105.0% -- -- -- --
iA (ours) 37.8% 44.9% 64.8% 33.0% 51.5% 63.9%
[…]
Table 4 summarizes the path planning performance on the Maze dataset, which is believed to be the most difficult map type due to its complicated routes. It can be seen that the FS+PPM algorithm performs poorly, even increasing 105% operation time and 8.9% search area than the classical A search. Additionally, iA shows the most significant runtime efficiency improvements on large maps, i.e., a 64.2% higher performance in Rt than Neural A. This further demonstrates the generalization ability of the IL framework.
@@ 5.1.5 Qualitative Performance
[…]
@@ 5.1.6 Baselines
[…]
@@
5.1.7 Platforms
To thoroughly evaluate the performance of local planning algorithms, we select four distinct simulated environments set up by (Cao et al., 2022), including indoor, garage, forest, and Matterport3D (Chang et al., 2017). In each of the environments, 30 pairs of start and goal positions are randomly selected in traversable areas for a comprehensive evaluation. It is worth noting that none of these environments was observed by iPlanner and the baseline data-driven methods during training. Specifically, a laptop with a 2.6 G Hz  Intel i9 CPU and an NVIDIA RTX 3080 GPU are used for running all methods.
5.1.8 Metrics
The success rate weighted by path length (SPL) will be used for evaluation (Anderson et al., 2018). It has been widely adopted by various path planning systems (Yokoyama et al., 2021; Ehsani et al., 2024).
5.1.9 Quantitative Performance
5.1.10 Efficiency Analysis
We evaluate iPlanner’s efficiency in Matterport3D. The robot follows 20 manual waypoints for global guidance but autonomously plans feasible paths and avoids local obstacles. As shown in Figure LABEL:fig:computing-time, iPlanner achieves an average 3×3$\times$3 × speedup over MPP. The executed trajectory is shown in Figure LABEL:fig:trajectory.
5.1.11 Real-world Tests
We adopt an ANYmal-legged robot (Hutter et al., 2016) which is equipped with an NVIDIA Jetson Orin and Intel Realsense D435 front-facing depth camera. The real-world experiments involve both dynamic and static obstacles in indoor, outdoor, and mixed settings. The robot is given a sequence of waypoints by a human operator, as illustrated in Figure 11. Specifically, the robot goes through areas including passing through doorways, circumventing both static and dynamic obstacles, and ascending and descending stairs. These trials are designed to test the adaptability and reliability of iPlanner under varied and unpredictable environmental conditions. During the test, iPlanner can generate feasible paths online and guide the robot through all areas safely and smoothly, which demonstrates the generalization ability of our IL framework.
@@ 5.2.1 Datasets
The VAP task in LogiCity (Li et al., 2024) integrates concept grounding and logical reasoning involving agent concepts and spatial relationships. It is particularly more challenging than other logical reasoning benchmarks with structured low-dimensional data, such as BlocksWorld (Dong et al., 2019) or knowledge graphs (Cropper and Morel, 2021). The main reasons are that LogiCity requires learning abstract concepts from varying agent distributions in an urban setting; additionally, it also involves high-dimensional image data with diverse visual appearances, however, as a comparison, BlocksWorld only contains binary vectors.
@@ 5.2.3 Baselines & Implementation Details
Since very little literature has studied FOL reasoning on RGB images, we don’t have many baselines. As a result, we adopt the method proposed in LogiCity (Li et al., 2024) that involves the SOTA visual encoder and reasoning engines. Particularly, we conduct experiments with two methods: (1) ResNet+NLM: ResNet as a visual encoder and NLM as the reasoning engine; (2) iLogic (ours): We train the “ResNet+NLM” within the IL framework, where the NLM is the LL problem. Specifically, the visual encoder includes a concept predictor and a relationship predictor. The concept predictor involves a ResNet FPN and an ROI-Align layer, followed by a 4-layer MLP to predict agent concepts (unary predicates). The relationship predictor includes a 3-layer MLP to encode paired agent information into agent relationships (binary predicates). The reasoning engine involves an NLM with breath as 3 and depth as 4. Two AdamW optimizers with a learning rate of 0.001 are adopted for the visual encoder and the reasoning engine, respectively.
Table 6: Evaluation results of the Easy mode in LogiCity.
[…]
Action Slow Normal Stop Overall
ResNet+NLM 0.671 0.644 0.995 0.724 0.00128
iLogic (ours) 0.678 0.679 0.995 0.740 0.00088
@@ 5.2.5 Efficiency Analysis
Table 7: Evaluation results of the Hard mode in LogiCity.
[…]
Action Slow Normal Fast Stop Overall
ResNet+NLM 0.552 0.575 0.265 0.999 0.400 0.00252
iLogic (ours) 0.459 0.598 0.338 0.999 0.442 0.00083
5.3 Optimal Control
5.3.1 Baselines
To demonstrate the effectiveness of the IL framework for jointly optimizing the data-driven IMU model and MPC, we select four baselines including (1) IMU+MPC: classic IMU integrator with a regular Diff-MPC; (2) IMU++MPC: data-driven IMU model with a regular Diff-MPC; (3) IMU+MPC+: classic IMU integrator with a Diff-MPC with learnable moment of inertia (MOI); and (4) iMPC (ours): IMU+ with MPC+ trained with IL framework. Specifically, “IMU+” refers to the data-driven IMU denoising and integration model from (Qiu et al., 2024). Additionally, for the wind disturbance test, we select an RL approach with a commonly used approach, proximal policy optimization (PPO) (Schulman et al., 2017) as the baseline.
5.3.2 Implementation Details
Due to efficiency considerations, we use a lightweight network for the data-driven IMU model, which includes a 2-layer multi-layer perceptron (MLP) as the encoder to get the feature embeddings from both the accelerometer and the gyroscope. Two decoders are implemented each comprising two linear layers to address the specific corrections needed for the accelerometer and the gyroscope. These decoders are dedicated to accurately predicting the necessary adjustments for each sensor type, thereby enhancing the overall precision of the IMU readings in dynamic conditions. For both the encoder and the decoder, we use GELU as the activation function for its effectiveness in facilitating smoother non-linear transformations. For RL, we take pre-integrated IMU measurements as the observation, output UAV action (thrust and torques) in OpenAI gym (Brockman et al., 2016), and designed the reward to minimize the difference between the current and target UAV pose.
Table 8: The performance of UAV attitude control under different initial conditions. The “IMU” means the RMSE (unit: ×$\times 10^{-3}\rm{rad}$) of attitude estimation for the corresponding method.
@@ 5.3.3 Simulation Environment
To accurately measure the performance of different controllers, we build a simulation environment for a quadrotor UAV with a standard 6 degree-of-freedom (DoF) dynamics model running at 1 K Hz  (Gabriel et al., 2007; Abdellah and Abdelaziz, 2004) and an onboard IMU sensor running at 200 Hz . To simulate real-world effects, such as environmental disturbance, actuator uncertainty, sensor noise, and other unknown behaviors, we inject zero-mean Gaussian noise with a standard deviation of 1e-4 to the control input and zero-mean Gaussian noise with a standard deviation of 8.73e-2 to the UAV attitude at 1K Hz . Additionally, we employ the sensor noise models documented by the Epson G365 IMU, including initial bias, bias instability, and random walk for both the accelerometer and gyroscope. This produces the same noise level as a typical UAV tracking system using a visual-inertial odometry system (Xu et al., 2023) for attitude estimation.
@@ 5.3.4 Metrics
To evaluate the performance of a controller, we use three widely-used metrics, including settling time (ST), root-mean-square error (RMSE), and steady-state error (SSE). Specifically, ST is the time for the UAV to enter and remain within ±$\pm 1.5\mathrm{\SIUnitSymbolDegree}$ of its final steady attitude, measuring how quickly the system settles; RMSE is the root-mean-square of the difference between estimated and the desired attitude; and SSE is the absolute difference between the steady and desired attitude, evaluating the control accuracy.
@@ 5.3.5 Robustness to Initial Conditions
Initial Offset Initial Error Initial Error Initial Error
50% $10\mathrm{\SIUnitSymbolDegree}$ $2.67\%$ $15\mathrm{\SIUnitSymbolDegree}$ $3.41\%$ $20\mathrm{\SIUnitSymbolDegree}$ $2.22\%$
5.3.6 Robustness to Wind Disturbance
We next evaluate the control performance under the wind disturbance to validate the robustness of the proposed approach. Specifically, we inject the wind disturbance by applying an external force $F=\frac{1}{2}\cdot C_{d}\cdot\rho\cdot A\cdot V^{2}$ and torque $\tau=r\times F$ to the nominal system based on the drag equation, where $C_{d}$ is the drag coefficient, $\rho$ is the air density, $A$ is the frontal area of the object facing the wind, $V$ is the wind speed, $r$ is the length of lever arm. Hence, for a fixed UAV configuration, different wind speeds $V$ lead to different disturbance levels. We conduct two sets of experiments: in the first set, we add impulse wind gust to the UAV in the direction of opposite heading during the hover with varying wind speed $10/15/20/45\ \mathrm{m}/\mathrm{s}$; in the second set, a step wind is added to the system in the same direction during hover and lasts for 0.3 s  for different wind speed $10/15/20/35\ \mathrm{m}/\mathrm{s}$. The external wind starts at 0.2 s  after entering a steady hover. During RL (PPO) training, we apply external wind disturbances of up to $10\mathrm{m}/\mathrm{s}$. In contrast, our hybrid approach with MPC provides robustness without requiring exposure to a wide range of wind conditions during training.
@@ 5.3.6 Robustness to Wind Disturbance
5.3.7 Efficiency Analysis
5.4 SLAM
5.4.1 Datasets
To ensure thorough evaluation, we adopt two widely-used datasets, including KITTI (Geiger et al., 2013) and EuRoC (Burri et al., 2016). They have diverse environments and motion patterns: KITTI incorporates high-speed long-range movements in driving scenarios, while EuRoC features aggressive motions in indoor environments.

### v7

@@ 5.1 Path Planning
In this section, we will conduct the experiments for iA* and iPlanner for global and local planning, respectively.
@@
To demonstrate the effectiveness of IL in global planning, we conduct both qualitative and quantitative evaluations for iA*.
@@ 5.1.1 Baselines
To evaluate the generalization ability of the proposed iA* framework, we select both classic and data-driven methods as the baselines. Specifically, we will compare iA* with the popular classic method, weighted A (pohl1970heuristic) and the latest imitation learning-based methods including neural A (yonetani2021nastar), focal search with path probability map (FS + PPM) and weighted A with correction factor (WA + CF) (kirilenko2023transpath).
@@ 5.1.2 Datasets
To evaluate iA* in diverse environments, we select three widely used datasets in path planning, Motion Planning (MP) (bhardwaj2017learning), Matterport3D (chang2017matterport3d), and Maze (maze-dataset). Specifically, the MP dataset contains a set of binary obstacle maps from eight environments; the Matterport3D dataset provides diverse indoor truncated signed distance function (TSDF) maps; and the Maze dataset contains sophisticated mazes-type maps. To better evaluate the generalization ability, we randomly sample the start and goal positions following (yonetani2021nastar) to provide more diverse maps. It is essential to note that all methods only use the MP dataset for training, and use Matterport3D and Maze for testing, which is to demonstrate their generalization abilities.
Table 2: Path planning evaluation on the MP dataset. The symbol ‘-’ denotes the method cannot provide meaningful results with that map size due to the large search space.
@@
WA+CF 47.3% 45.3% - - - -
FS+PPM -8.9% -105.0% - - - -
iA* (ours) 37.8% 44.9% 64.8% 33.0% 51.5% 63.9%
5.1.3 Metrics
To evaluate the performance of different algorithms, we adopt the widely used metrics including the reduction ratio of node explorations (Exp) (yonetani2021nastar) and the reduction ratio of time (Rt). Exp is the ratio of the reduced search area relative to the classical A, which can be defined as $\textit{Exp}=100\times\frac{S^{*}-S}{S^{*}}$, where $S$ represents the search area of the method and $S^{*}$ is the search area of classical A. Rt is the ratio of saved time relative to the classical A, which can be defined as $\textit{Rt}=100\times\frac{t^{*}-t}{t^{*}}$, where $t$ is the runtime of the method and $t^{*}$ is the runtime of classical A. Intuitively, Exp aims to measure the search efficiency while Rt aims to measure the runtime efficiency.
5.1.4 Quantitative Performance
Table 2 summarizes the performance on the MP dataset. It can be seen that although iA* ranked 2nd for map size of $64\times 64$, it achieves outstanding performance for larger maps with sizes $128\times 128$ and $256\times 256$. Specifically, it improves the runtime efficiency Rt by 23% for map size $256\times 256$ compared to Neural A, the SOTA method. This indicates that iA* has an advantage in larger maps, demonstrating the effectiveness of the upper-level predicted cost function for reducing the search space for LL path planning. We observe that FF+PPM and WA+CF are limited to a fixed shape of input instances due to their integration of the attention mechanism. Their network architectures include fully connected (FC) layers, which require predefined input and output shapes. The number of parameters within these FC layers varies with different input and output shapes. Consequently, adjustments and re-training procedures are necessary to accommodate varying input instance shapes.
Table 3 summarizes the path planning performance on the Matterport3D dataset, where iA* outperforms all methods in all metrics and map sizes. Specifically, iA* achieves a 12.7% and a 37.4% higher performance in search efficiency Exp and runtime efficiency Rt than the SOTA method neural A, respectively. It is worth noting that FS+PPM, the 1st ranked method on the $64\times 64$ map category in the MP dataset, achieves the worst performance on the Matterport3D dataset, indicating that FS+PPM is fine-tuned overfitting to the MP dataset and cannot generalize to other types of maps.
Table 4 summarizes the path planning performance on the Maze dataset, which is believed to be the most difficult map type due to its complicated routes. It can be seen that the FS+PPM algorithm performs poorly, even increasing 105% operation time and 8.9% search area than the classical A search. Additionally, iA* shows the most significant runtime efficiency improvements on large maps, i.e., a 64.2% higher performance in Rt than Neural A. This further demonstrates the generalization ability of the IL framework.
@@ 5.1.5 Qualitative Performance
For the MP dataset, as shown in the 1st row of Figure 9, it is observed that all the methods have a smaller search space than the classical A and our iA* has the smallest search space, indicating that iA* achieves the best overall efficiency. It is also observed that the path found by Neural A is not smooth and all the other data-driven methods show slight differences in terms of path shape and length. This further indicates the effectiveness of our iA* framework.
The 2nd row of Figure 9 displays the path planning results from the Matterport3D dataset, which mainly focuses on indoor scenarios. It can be seen that the search space of FS+PPM and WA+CF are larger than the other methods. It is worth noting that FS+PPM has a larger search space than classical A. Besides, all the optimal paths found by different methods are close to the optimal path of classical A.
In the 3rd row of Figure 9, we show the performance of all baseline methods for the maze-type maps from the Maze dataset. From these examples, it can be observed that all data-driven methods successfully improve the search efficiency than classical A. Additionally, iA* has the smallest search space, which means that it can even greatly improve the search efficiency of maze-level complicated maps.
@@ 5.1.6 Baselines
To evaluate the robustness and efficiency of IL in local path planning, we will compare iPlanner against both classic and data-driven methods. Specifically, we will compare with one of the SOTA classic planners, motion primitives planner (MPP) (zhang2020falco; cao2022autonomous), which received the 1st place for most explored sectors in the DARPA Subterranean (SubT) challenge (darpa_subt). Additionally, we will also compare with data-driven methods, including the supervised learning (SL)-based method (loquercio2021learning) and the reinforcement learning (RL)-based method (hoeller2021learning), which will be simply denoted as SL and RL methods, respectively.
@@
5.1.7 Platforms
To thoroughly evaluate the performance of local planning algorithms, we select four distinct simulated environments set up by (cao2022autonomous), including indoor, garage, forest, and Matterport3D (chang2017matterport3d). In each of the environments, 30 pairs of start and goal positions are randomly selected in traversable areas for a comprehensive evaluation. It is worth noting that none of these environments was observed by iPlanner and the baseline data-driven methods during training. Specifically, a laptop with a 2.6 G Hz  Intel i9 CPU and an NVIDIA RTX 3080 GPU are used for running all methods.
5.1.8 Metrics
The success rate weighted by path length (SPL) will be used for evaluation (anderson2018evaluation). It has been widely adopted by various path planning systems (yokoyama2021success; ehsani2024spoc).
5.1.9 Quantitative Performance
5.1.10 Efficiency Analysis
We evaluate iPlanner’s efficiency in Matterport3D. The robot follows 20 manual waypoints for global guidance but autonomously plans feasible paths and avoids local obstacles. As shown in Figure LABEL:fig:computing-time, iPlanner achieves an average $3\times$ speedup over MPP. The executed trajectory is shown in Figure LABEL:fig:trajectory.
5.1.11 Real-world Tests
We adopt an ANYmal-legged robot (hutter2016anymal) which is equipped with an NVIDIA Jetson Orin and Intel Realsense D435 front-facing depth camera. The real-world experiments involve both dynamic and static obstacles in indoor, outdoor, and mixed settings. The robot is given a sequence of waypoints by a human operator, as illustrated in Figure 11. Specifically, the robot goes through areas including passing through doorways, circumventing both static and dynamic obstacles, and ascending and descending stairs. These trials are designed to test the adaptability and reliability of iPlanner under varied and unpredictable environmental conditions. During the test, iPlanner can generate feasible paths online and guide the robot through all areas safely and smoothly, which demonstrates the generalization ability of our IL framework.
@@ 5.2.1 Datasets
The VAP task in LogiCity (li2024logicity) integrates concept grounding and logical reasoning involving agent concepts and spatial relationships. It is particularly more challenging than other logical reasoning benchmarks with structured low-dimensional data, such as BlocksWorld (dong2019nlm) or knowledge graphs (cropper2021popper). The main reasons are that LogiCity requires learning abstract concepts from varying agent distributions in an urban setting; additionally, it also involves high-dimensional image data with diverse visual appearances, however, as a comparison, BlocksWorld only contains binary vectors.
Table 6: Evaluation results of the Easy mode in LogiCity.
Metric Recall Rate wAcc $\textrm{Var}(\text{wAcc})$
Action Slow Normal Stop Overall
ResNet+NLM 0.671 0.644 0.995 0.724 0.00128
iLogic (ours) 0.678 0.679 0.995 0.740 0.00088
@@ 5.2.3 Baselines & Implementation Details
Since very little literature has studied FOL reasoning on RGB images, we don’t have many baselines. As a result, we adopt the method proposed in LogiCity (li2024logicity) that involves the SOTA visual encoder and reasoning engines. Particularly, we conduct experiments with two methods: (1) ResNet+NLM: ResNet as a visual encoder and NLM as the reasoning engine; (2) iLogic (ours): We train the “ResNet+NLM” within the IL framework, where the NLM is the LL problem. Specifically, the visual encoder includes a concept predictor and a relationship predictor. The concept predictor involves a ResNet FPN and an ROI-Align layer, followed by a 4-layer MLP to predict agent concepts (unary predicates). The relationship predictor includes a 3-layer MLP to encode paired agent information into agent relationships (binary predicates). The reasoning engine involves an NLM with breath as 3 and depth as 4. Two AdamW optimizers with a learning rate of 0.001 are adopted for the visual encoder and the reasoning engine, respectively.
Table 7: Evaluation results of the Hard mode in LogiCity.
Metric Recall Rate wAcc $\textrm{Var}(\text{wAcc})$
Action Slow Normal Fast Stop Overall
ResNet+NLM 0.552 0.575 0.265 0.999 0.400 0.00252
iLogic (ours) 0.459 0.598 0.338 0.999 0.442 0.00083
@@ 5.2.5 Efficiency Analysis
Table 8: The performance of UAV attitude control under different initial conditions. The “IMU” means the RMSE (unit: $\times 10^{-3}\rm{rad}$) of attitude estimation for the corresponding method.
@@
5.3 Optimal Control
5.3.1 Baselines
To demonstrate the effectiveness of the IL framework for jointly optimizing the data-driven IMU model and MPC, we select four baselines including (1) IMU+MPC: classic IMU integrator with a regular Diff-MPC; (2) IMU++MPC: data-driven IMU model with a regular Diff-MPC; (3) IMU+MPC+: classic IMU integrator with a Diff-MPC with learnable moment of inertia (MOI); and (4) iMPC (ours): IMU+ with MPC+ trained with IL framework. Specifically, “IMU+” refers to the data-driven IMU denoising and integration model from (qiu2023airimu). Additionally, for the wind disturbance test, we select an RL approach with a commonly used approach, proximal policy optimization (PPO) (schulman2017proximal) as the baseline.
5.3.2 Implementation Details
Due to efficiency considerations, we use a lightweight network for the data-driven IMU model, which includes a 2-layer multi-layer perceptron (MLP) as the encoder to get the feature embeddings from both the accelerometer and the gyroscope. Two decoders are implemented each comprising two linear layers to address the specific corrections needed for the accelerometer and the gyroscope. These decoders are dedicated to accurately predicting the necessary adjustments for each sensor type, thereby enhancing the overall precision of the IMU readings in dynamic conditions. For both the encoder and the decoder, we use GELU as the activation function for its effectiveness in facilitating smoother non-linear transformations. For RL, we take pre-integrated IMU measurements as the observation, output UAV action (thrust and torques) in OpenAI gym (brockman2016openai), and designed the reward to minimize the difference between the current and target UAV pose.
@@ 5.3.3 Simulation Environment
To accurately measure the performance of different controllers, we build a simulation environment for a quadrotor UAV with a standard 6 degree-of-freedom (DoF) dynamics model running at 1 K Hz  (gabriel2007quadrotor; abdellah2004dynamic) and an onboard IMU sensor running at 200 Hz . To simulate real-world effects, such as environmental disturbance, actuator uncertainty, sensor noise, and other unknown behaviors, we inject zero-mean Gaussian noise with a standard deviation of 1e-4 to the control input and zero-mean Gaussian noise with a standard deviation of 8.73e-2 to the UAV attitude at 1K Hz . Additionally, we employ the sensor noise models documented by the epsonG365IMU IMU, including initial bias, bias instability, and random walk for both the accelerometer and gyroscope. This produces the same noise level as a typical UAV tracking system using a visual-inertial odometry system (xu2023airvo) for attitude estimation.
Initial Offset Initial Error Initial Error Initial Error
50% $10\mathrm{\SIUnitSymbolDegree}$ $2.67\%$ $15\mathrm{\SIUnitSymbolDegree}$ $3.41\%$ $0\mathrm{\SIUnitSymbolDegree} 2.2$$22\%$
@@ 5.3.4 Metrics
To evaluate the performance of a controller, we use three widely-used metrics, including settling time (ST), root-mean-square error (RMSE), and steady-state error (SSE). Specifically, ST is the time for the UAV to enter and remain within $\pm 1.5\mathrm{\SIUnitSymbolDegree}$ of its final steady attitude, measuring how quickly the system settles; RMSE is the root-mean-square of the difference between estimated and the desired attitude; and SSE is the absolute difference between the steady and desired attitude, evaluating the control accuracy.
@@ 5.3.6 Robustness to Wind Disturbance
5.3.7 Efficiency Analysis
5.4 SLAM
5.4.1 Datasets
To ensure thorough evaluation, we adopt two widely-used datasets, including KITTI (geiger2013vision) and EuRoC (burri2016euroc). They have diverse environments and motion patterns: KITTI incorporates high-speed long-range movements in driving scenarios, while EuRoC features aggressive motions in indoor environments.